OpenStudy (anonymous):

secx + tan=1 are you allowed to square these ?

OpenStudy (anonymous):

yeah you are allowed. Will you post the problem in which you are going to apply ?

OpenStudy (anonymous):

allowed

OpenStudy (anonymous):

It says solve on the interval 0<=x<2pi secx+tanx=1 I got to -1+1/cos+sin/cos=0

OpenStudy (anonymous):

u know the answer

OpenStudy (anonymous):

will i post a solution in a non-trivial way?.@Lorayne

OpenStudy (anonymous):

No I am looking at the examples and lost... no idea why you are allowed to square both sides to make it sinx^2+Cosx^2=1

OpenStudy (anonymous):

@Lorayne check the attachment. I am sure you can proceed after that

OpenStudy (anonymous):

Thank you I am looking at it now :)

OpenStudy (anonymous):

@Lorayne did you get the solution? Next all you have to do is just find the x values in the given range.

OpenStudy (anonymous):

Sorry this is my first one of these and still learning how you got the sin on the bottom... did you use sin^2+1 to exchange it ?

OpenStudy (anonymous):

cos2x =cos ^2 x- sin ^2 x. I used this formula

OpenStudy (anonymous):

ok, looking.

OpenStudy (anonymous):

ok how do you get the pi over 4 ?

OpenStudy (anonymous):

tan (a+b)=(tan a+tan b)/(1-tan a tanb) : formula.

OpenStudy (anonymous):

ok I just learned that in the last section. I don't have it in my example so I wasn't sure

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

and when tan A=tan B, then A=npi+B, where n is an integer

OpenStudy (anonymous):

I am going back in the previous section and rereading lol I am missing steps